1.0 mole of a gas occupies a volume of 22.4 L gas at 0 ̊C and 1 atm.

What would be the volume of 7.5 mol of the gas at the same temperature

and pressure?


1
Expert's answer
2021-09-28T01:04:15-0400

"\frac{P1V1}{n1T1}=\frac{P2V2}{n2T2}"

"\frac{\times{1}{22.4}}{\times{1}{273}}=\frac{\times{1}{x}}{\times{7.5}{273}}"

"0.08= \frac{x}{2047.5}"

"x = 163.8 L"